The allowance for uncollectible accounts had a credit balance of $10,000 at the beginning of 2021.

Johnson Company calculates its allowance for uncollectible accounts as 10% of its ending balance in gross accounts receivable. The allowance for uncollectible accounts had a credit balance of $10,000 at the beginning of 2021. No previously written-off accounts receivable were reinstated during 2021. At 12/31/2021, gross accounts receivable totaled $166,700, and prior to recording the adjusting entry to recognize bad debts expense for 2021, the allowance for uncollectible accounts had a debit balance of 18,300. 


Required:

 1. What was the balance in gross accounts receivable as of 12/31/2020?

 2. What journal entry should Johnson record to recognize bad debt expense for 2021?

 3. Assume Johnson made no other adjustment of the allowance for uncollectible accounts during 2021. Determine the amount of accounts receivable written off during 2021.

 4. If Johnson instead used the direct write-off method, what would bad debt expense be for 2021? 
Answer #1

Solution 1:

Gross accounts Receivable at 12/31/2020 = Allowance Account balance at beginning / 10%

= $10,000 / 10% = $100,000

Solution 2:

Year Account Title Debit Credit
2021 Bad debt expense Dr ($166700*10% + 18300) $34,970
    To Allowance for Doubtful Accounts $34,970

Solution 3:

Accounts receivable written off = Beginning balance of Allowance Account - Ending Balance of Allowance account

= $10000 - (- $18300) = $10000 + $18300 = $28,300

Solution 4:

Bad debt expense for 2021 (direct write off method) = Amount written off = $28,300
Answer #2

Solution 1:

Gross accounts receivable as of 12/31/2020 = Allowance for uncollectible account balance on January 1, 2021/ 10% = $10000/ 10% = $100,000

Solution 2:

Year General journal Debit Credit
2021 Bad debt expense ($166700*10% + $18300) 34970
Allowance for uncollectible account   34970

Solution 3:

Accounts receivable written off = $10000 - (-$18300) = $28,300

Solution 4:

Bad debt expense under direct write off method = $28,300